OpenStudy (anonymous):

Polygon ABCD is translated to create polygon A′B′C′D′. Point A is located at (1, 5), and point A′ is located at (-2, 3). Which expression defines the transformation of any point (x, y) to (x′, y′) on the polygons?

OpenStudy (godlovesme):

options?

OpenStudy (anonymous):

@Godlovesme x′ = x − 3 y′ = y − 2 x′ = x − 2 y′ = y − 3 x′ = x − 1 y′ = y − 8 x = x′ + 3 y = y′ + 2

OpenStudy (anonymous):

I figured it out, it's A